Smithson Company uses a job-order costing system and has two manufacturing departments—Molding and Fabrication. The company provided the following estimates at the beginning of the year: Molding Fabrication Total Machine-hours 23,000 33,000 56,000 Fixed manufacturing overhead costs $ 730,000 $ 220,000 $ 950,000 Variable manufacturing overhead per machine-hour $ 5.00 $ 5.00 During the year, the company had no beginning or ending inventories and it started, completed, and sold only two jobs—Job D-75 and Job C-100. It provided the following information related to those two jobs: Job D-75: Molding Fabrication Total Direct materials cost $ 377,000 $ 329,000 $ 706,000 Direct labor cost $ 210,000 $ 170,000 $ 380,000 Machine-hours 18,000 5,000 23,000 Job C-100: Molding Fabrication Total Direct materials cost $ 280,000 $ 290,000 $ 570,000 Direct labor cost $ 100,000 $ 220,000 $ 320,000 Machine-hours 5,000 28,000 33,000 Smithson had no overapplied or underapplied manufacturing overhead during the year. 3.value: 1.66 pointsRequired information Assume Smithson uses a plantwide overhead rate based on machine-hours. Required: 1-a. Compute the predetermined plantwide overhead rate. (Round your answer to 2 decimal places.) 1-b. Compute the total manufacturing costs assigned to Job D-75 and Job C-100. (Round your intermediate calculations to 2 decimal places.) 1-c. If Smithson establishes bid prices that are 120% of total manufacturing costs, what bid price would it have established for Job D-75 and Job C-100? (Round your intermediate calculations to 2 decimal places.) 1-d. What is Smithson’s cost of goods sold for the year? (Round your intermediate calculations to 2 decimal places.)
Answer #1

Answer

  • All working forms part of the answer
  • Requirement 1 ‘a’

Working

Molding Department:

Fabrication Department:

Total

A

Variable Overhead rate

$                 5.00

$                        5.00

B

Total machine hours

                23,000

                       33,000

                   56,000

C = A x B

Total Variable Overhead

$          115,000

$                 165,000

$             280,000

D

Total Fixed Overheads

$          730,000

$                 220,000

$             950,000

E = C+D

Total manufacturing Overhead

$          1,230,000

F

Total machine hours

                   56,000

G = E/F

Plant Wide Overhead rate

$                  21.96 = Answer

  • Requirement 1 ‘b’

Job D 75

Job C 100

Molding Department:

Direct material cost

$          377,000

$                 280,000

Direct Labor cost

$          210,000

$                 100,000

Manufacturing Overhead

$          395,280

$                 109,800

Fabrication Department:

Direct material cost

$          329,000

$                 290,000

Direct Labor cost

$          170,000

$                 220,000

Manufacturing Overhead

$          109,800

$                 614,880

Total manufacturing cost assigned

$      1,591,080 = Answer

$             1,614,680 = Answer

  • Requirement 1 ‘c’

Job D 75

Job C 100

A

Total manufacturing cost assigned

$                              1,591,080

$     1,614,680

B = A x 120%

Bid Price

$                              1,909,296

$     1,937,616

  • Requirement 1 ‘d’

Cost of Goods Sold = 1591080 + 1614680 = $ 3,205,760